a children’s liquid cold medicine has a density of 1.23 g/mL. if a child is to take 2.5 tsp in a dose, what is the mass in grams of this dose?
Answer #1

We know that, density=mass×volume

Or,. d=M×V

So in the question it is given that, d=1.23 g/ml. And also volume is given equal to 2.5 tablespoon.

Also we know that, 1tsp = 15 ml. Hence 2.5 tsp has 2.5×15 ml.

So volume will be 2.5×15

By putting the given values we can find out the mass.

M=d×V

=1.23×2.5×15

=46.125g

Hence, mass in the dose is 46.125g